In the staff room there is 7/8 of a cake left over from a meeting.
If 14 members of staff would all like a piece, what fraction will they each receive?

Answer by ptaylor(2198)

Let x represent the fraction that each will get
Now we know that that when we add all the 14 fractions up they must equal 7/8.
Soooo our equation to solve is:
14*x=7/8
x=(7/8)/14 multiply numerator and denominator by 1/14 to get rid of the complex fraction.
x=(7/8)(1/14)/1=1/16 fraction that each will get
CK
(1/16)*14=7/8
7/8=7/8
Hope this helps---ptaylor